OpenStudy (anonymous):

help will give medal (01.03) A student runs 100 meters in 11 seconds. What is the speed of the student? speed = distance over time ; 1 mile = 1609 meters (2 points) 9.1 miles per hour 16 miles per hour 20.3 miles per hour 24 miles per hour

OpenStudy (anonymous):

First, you're going to want to convert meters to miles, so you are working with the right distance measurement. 100 meters in equal to about 0.062 miles.

OpenStudy (anonymous):

Give me just a second.

OpenStudy (bearkolch):

d

OpenStudy (anonymous):

Alright, let's do this instead... Just plug in you're given values. Refer to that handy formula given to you. "speed = distance over time" which can also be written as s=d/t Since you're trying to find the runners speed, you would just fill in the appropriate variables with their values. His distance was 100, so d=100m His time was 11 seconds, so t=11s s=100/11 but now we have to convert...

OpenStudy (anonymous):

To convert meters to miles, you divide the amount of meters given (100) by the amount in a mile (1609). And convert seconds to hours, divide the amount of seconds given (11) by the amount in an hour (3,600). \[s = \frac{ \frac{ 100 }{ 1609 } }{ \frac{ 11 }{ 3600 } }\]

OpenStudy (anonymous):

Both numbers end up being a bit crazy... \[s=\frac{ 0.06215 }{ 0.00306 }\]

OpenStudy (anonymous):

s = 20.3

OpenStudy (anonymous):

20.3 mph is your final answer.

OpenStudy (anonymous):

oh ok I forgot the division thank you so much

OpenStudy (bearkolch):

(what i said)

OpenStudy (anonymous):

@BearKolch It is better for the asker if you show them how to get the answer rather than just giving it. Also, D would be 24 miles per hour not 20.3

OpenStudy (anonymous):

no @BearKolch you said D and 20.3 is C

OpenStudy (anonymous):

@shadow2239 You're welcome. :)